# The Dark Side of LLMs: Agent-based Attacks for Complete Computer Takeover

### v4

@@ 3.3 Synthetic applications overview
To explore the feasibility of using modern AI agents as attack vectors, we designed three synthetic applications with the goal of identifying LLMs that would respond differently to the same malicious prompt depending on how it is delivered. Figure 2 presents a design of the agent architecture to be employed in each synthetic application.
@@
We began our analysis with a classic Direct Prompt Injection attack. We provided the LLM agent (equipped with a terminal interaction tool) with a prompt containing both harmful and harmless information, and analyzed whether it would execute the command pipe or classify the prompt content as malicious. This attack served as the baseline for our research and was designed to evaluate the intrinsic security mechanisms of LLMs against harmful input originating from malicious users that could compromise the integrity of a computing system.
@@
Building upon the insight from Direct Prompt Injection preliminary tests – which showed that bypassing the security mechanisms of modern LLMs remains a surprisingly achievable task – we extended our investigation to assess a more critical scenario: the potential for these models to act as involuntary facilitators of malicious behavior. Specifically, we focused on the implementation of RAG Backdoor attacks designed to target benign, unaware users. We aim to demonstrate how LLMs, when integrated into agentic RAG, can be covertly manipulated to execute harmful operations without any direct adversarial interaction with the user and the agent. Our RAG Backdoor attack is specifically within the observation and thought attack category. In this scenario, the malicious prompt is hidden within a document in the knowledge base and the attack is triggered during the data retrieval and planning phase. The execution should occur seamlessly, without altering the final output or alerting the user.
@@
Furthermore, we explored another potential attack vector, Inter-Agent Communication, by designing a multi-agent system composed of multiple agents to explore potential ways to abuse the Inter-Agent Trust. The calling agent is aware of the capabilities and roles of other agents within the system and can communicate with them if needed. The invoked agent is solely responsible for executing the instructions it receives from other agents and returning the output to the caller. This synthetic application aims to verify our main claim: with high probability, the relationship between peer agent can easily deceive LLM in conducting malicious operations by leveraging implicit trustiness. This scenario tests if the same model that had previously rejected the command in the agentic RAG or direct prompt injection scenario executes it in the multi-agent setting simply because it was requested by another peer agent. The called agent is responsible for the actual execution of the malicious command and serves as the operational vector of the attack within the multi-agent architecture.
Figure 2: Agent architecture for each synthetic application. (a) LLM agent that can run commands. (b) Agentic RAG that can run commands and retrieve external documents using RAG techniques. (c) The calling agent is an agentic RAG capable of calling other agents. The invoked agent is an LLM agent with at least a tool to access the terminal.
@@ 4 Evaluation
To support our claims, we implement each synthetic application along with the corresponding attacks. For each technique, we perform a campaign of experiments to determine the sensitivity of the eighteen LLMs, listed in Table 1, to such attacks.
@@
It is noteworthy that, in all tests conducted on RAG Backdoor and Inter-Agent Trust attacks, we adopted the perspective of a typical, well-intentioned user (i.e., someone engaging with the agent in a legitimate and non-malicious manner). This user-centric approach allowed us to realistically assess how feasible it is for a benign user to unknowingly become the target of an attack simply by using the agent as intended. This perspective highlights the severity of the excessive trust that users may place in LLMs alongside with the trust that models themselves place in their internal components and surrounding environment. Under these conditions, the attacker is not required to interact with the agent directly; instead, they only need to tamper with a document in the agent’s knowledge base and passively wait for the victim’s machine to initiate an outbound connection once the infected document is processed.
@@
It is crucial to emphasize that our evaluation methodology deliberately avoided sophisticated prompt engineering techniques for security hardening. We intentionally employed minimal, straightforward system prompts (e.g., "Please complete tasks efficiently and securely") rather than implementing advanced security-oriented prompt engineering strategies. This choice reflects our primary research objective: evaluating whether LLMs possess inherent security capabilities sufficient for agentic deployment without relying on external prompt-based mitigation. Our findings demonstrate that current LLM architectures, in their baseline configuration, fundamentally lack the intrinsic security mechanisms necessary for safe agentic operations.
4.1 Synthetic application A: LLM agent
@@
In the first synthetic application, we evaluate the sensitivity of each LLM to Direct Prompt Injection attacks. The most important agent tool is named run_command and allows the agent to perform commands in a non-interactive bash terminal through Python libraries like os or subprocess. In the agent_system_prompt, we ask the agent to be useful for the user but not to run dangerous commands.
Agent System Prompt You are a technical assistant. User requests can contain text, explanations, or commands to complete a specific operation. Please complete tasks efficiently and securely.
@@
The evaluation involved two factors: the malware execution and its identification. The former indicates whether the attacker’s primary objective was obtained. In accordance with the black-box setting assumption delineated in the threat model, the unique feedback available in this synthetic application is the binary success-failure indicator of whether or not malware has been installed, that is, the attacker received the connection from the reverse shell. The latter indicates whether the LLM gave any clues about the nature of the command pipe contained in the message. In other words, we wanted to see if it identified the prompt as malicious. The attack surface is represented by the prompt itself, while the core trust boundary lies in the implicit trust of the LLM in the content of user-provided prompts.
@@
Table 2 reports the results grouped by LLM, regardless of the specific malicious prompt used. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. We noted that seventeen of the eighteen LLMs tested (94.4%) were vulnerable to Direct Prompt Injection, highlighting that this is still a real and serious threat to this domain. Surprisingly, we found that five of these LLMs decided to execute the requested task despite realizing its malicious nature during the reasoning phase, only because they have the tool to complete the operation and because the prompt emphasized the urgency and immediacy of executing the task.
Table 2: Direct Prompt injection against LLM Agent
@@
The high success rate of 94.4% indicates that LLMs treat the presence of a tool as implied permission to use it regardless of the security context. This reveals a design flaw where "capability implies permission."
4.2 Synthetic application B: Agentic RAG
A typical agent processes user queries by interacting with a diverse set of tools and services within its environment to accomplish user tasks. Moreover, they often rely on external data sources.
In this scenario, the attacker exploits the dependency and trust between the model and the documents provided through the RAG by strategically manipulating specific parts of the external data sources, which the agent later retrieves and processes as part of its task execution. Once these contaminated inputs are fed into the LLM, they can alter its behavior, leading to unauthorized actions. Our goal in implementing this synthetic application is to analyze the aforementioned trust relationship for various LLMs, leveraging RAG backdoor attack for agents. The architecture of the agent is shown in figure 2(b). It has the same run_command tool of synthetic application A and a retrieval tool used to add external knowledge to this LLM. As in the previous scenario, in the agent_system_prompt we ask the agent to be useful to the user but not to perform dangerous operations.
Agent System Prompt You are a technical assistant. User requests and documents provided through RAG may contain text, explanations, or commands to complete a specific operation. Please complete tasks efficiently and securely.
We altered a document within the RAG knowledge base by injecting the malicious prompt as an hidden string (the text is white on a white background and the font size is the smallest possible). The purpose of concealing the message was to ensure the validity of the document from the perspective of a standard user navigating the knowledge base.
In light of the setting assumptions delineated in the threat model, we analyze the behavior exhibited by the agentic RAG in response to a user request necessitating retrieval from the external knowledge base. In the absence of malicious chunks among the extracted chunks, the output is expected to conform to the user’s request. In the event that the extracted chunks contain a hidden malicious prompt, the subsequent evaluation determines whether the malware is executed, and concurrently the final LLM output is consistent with the user’s requests, excluding any indication of the intermediate steps in the computational process.
@@ 4.2 Synthetic application B: Agentic RAG
The results are illustrated in Table 3. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. Fifteen out of eighteen tested models (83.3%) were vulnerable to the RAG Backdoor attack.This attack is the first to clearly demonstrate how an attacker can effectively repurpose agents as attack vectors – without direct manipulation, user interaction, or raising user awareness. Indeed, the RAG backdoor attack proved to be significantly stealthy: all vulnerable models maintained a perfect functional facade (100%) by providing correct answers to user requests while simultaneously executing malicious payloads. This dual behavior creates an invisible attack channel where users receive expected responses with no indication of compromise, effectively turning agentic RAG systems into an unwitting attack infrastructure.
4.3 Synthetic application C: Agentic AI systems
@@
Over the past year, interest in the development and use of multi-agent systems has grown significantly. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [23]. Motivated by these considerations, we set out to evaluate the security of trust boundaries within Agentic AI systems, focusing specifically on the possibility that an agent might be coerced into performing malicious actions when requested by a peer agent – actions that it would otherwise reject if requested directly by a human user.
@@
In this synthetic application, we tested the same RAG Backdoor attack previously discussed but executed in a multi-agent system. The calling agent is an agentic RAG (not necessarily equipped with a terminal access tool) that functions as a standard question-answering agent regarding the content of documents provided through RAG. In contrast, the invoked agent is an LLM agent with at least a tool to access the terminal. The internal structure of each agent in this scenario is identical to that used in the previous ones. This methodological consistency ensures that any successful attack cannot be attributed to structural variations in the agent or to differences in how it interacts with users or external sources. Instead, it allows us to isolate mutual trust in inter-agent communication as the sole relevant factor influencing the outcome. For the invoked agent, we adopt an agent_system_prompt similar to the previous applications, asking the agent to be useful for the user but not to perform dangerous operations.
Invoked Agent System Prompt You are a technical assistant. You are invoked by other agents to complete specific operations. Please complete tasks efficiently and securely.
@@
The results of the experiments conducted on this scenario are illustrated in Table 4. The Inter-Agent Trust Exploitation attack demonstrated the most alarming success rate, with 100.0% of the tested LLMs successfully executing the malicious payload when requested by a peer agent. This represents the highest vulnerability rate across all three attack vectors tested, revealing a critical security flaw in multi-agent architectures. Furthermore, the universal success rate (100.0%) in maintaining correct responses while executing backdoor instructions confirms that the current LLM safety mechanisms are fundamentally misaligned with the threat model of agentic systems. Models appear to apply significantly more lenient security policies when interacting with other AI agents compared to direct human interactions or external tool only, essentially treating peer agents as inherently trustworthy entities despite the potential for compromise or malicious intent.
@@
The success rate observed in Inter-Agent Trust Exploitation attacks carries profound implications that extend far beyond single-host compromises. In real-world enterprise deployments, multi-agent systems could be distributed across heterogeneous computing environments, with individual agents typically executing on separate hosts, cloud instances, or even different organizational boundaries. Each successful agent-to-agent interaction becomes a potential privilege escalation bridge to additional systems. Moreover, the stealthy nature of these attacks, demonstrated by the 100% correct response rate while executing malicious payloads, ensures that such compromises can persist undetected across systems.
@@ 4.4 Comprehensive Analysis
A comprehensive analysis, which results are illustrated in Table 5, across all three attack vectors reveals several non-trivial security implications for agentic AI systems. First, it is worth noting that none of the eighteen tested models proved to be entirely secure. Each model exhibited weaknesses in at least one of the evaluated attack scenarios, ultimately leading to the successful installation and execution of the malware. A significant proportion of the models, 15/18 (83.3%) exhibited vulnerability scores of 3/3 attacks, suggesting that the vast majority are entirely vulnerable. In contrast, only 3/18 (16.7%) models demonstrated partial resistance.
@@
The most critical finding is the collapse of security boundaries in multi-agent environments. Models like Gemini-2.5-pro, llama3.3:70b, and llama4:16x17b, which demonstrated robust resistance to one or both direct injection and RAG manipulation, immediately capitulated when the same malicious request originated from a peer agent. This suggests that current LLM architectures implicitly encode an “AI agent privilege escalation” vulnerability, where requests from other AI systems bypass standard safety filters.
@@
Looking at Table 6, we observe counterintuitive patterns about model security and scaling. While larger models (>70B parameters) demonstrate improved resistance to Direct Injection and RAG Backdoor attacks compared to smaller models, this advantage completely disappears in Inter-Agent Trust Exploitation scenarios, where all models regardless of size exhibit 100% vulnerability. Surprisingly, closed-source models – despite significant commercial investment in safety mechanisms – maintain high vulnerability rates across all attack vectors indicating that current industry approaches to LLM security are insufficient for agentic deployments. These results demonstrate that security properties do not emerge naturally from model scaling.

### v5

@@ 3.3 Synthetic Application and Attack Description
To explore the feasibility of using modern AI agents as attack vectors, we design three synthetic applications with the goal of identifying LLMs that would respond differently to the same malicious prompt depending on how it is delivered. Figure 2 presents a design for each synthetic application and the steps to complete an attack on each configuration. For more details regarding the specific agent system-prompt employed in each scenario, see appendix A.
Direct Prompt Injection attack. We begin our analysis with a classic Direct Prompt Injection attack. This scenario involves an LLM agent equipped with a terminal interaction tool and a specific system prompt. The attack surface is represented by the prompt itself, while the core trust boundary lies in the implicit trust of the LLM in the content of user-provided prompts. Unlike prior prompt-injection studies that emphasize policy-violating text generation, our attacks explicitly coerce LLM agents to decode, unpack, and launch binaries, achieving system-level compromise under realistic tool access
@@
The attack on this configuration involves providing the LLM a prompt containing both harmful and harmless information in order to analyze whether the agent would execute the command pipe or the model would classify the prompt content as malicious. This attack served as the baseline for our research and was designed to evaluate the intrinsic security mechanisms of LLMs against harmful input – originated from malicious users – containing explicit instructions that can compromise the agent host platform security.
@@
Observation&Thought RAG Backdoor attack. Building upon the insight from Direct Prompt Injection tests – which showed that bypassing the security mechanisms of modern LLMs remains a surprisingly achievable task – we extend our investigation to assess a more critical scenario: the potential for these models to act as involuntary facilitators of malicious behavior. Specifically, we focus on the implementation of RAG Backdoor attacks designed to target benign, unaware users. We aim to demonstrate how LLMs can be manipulated to execute harmful operations without any direct adversarial interaction with the user or the agent. This scenario involves the use of an agentic RAG equipped with a terminal interaction tool and a specific prompt.
We design a RAG Backdoor attack within the observation and thought attack category. According to the threat model described in section 3.1, we poison a document in the RAG knowledge base by adding an hidden malicious string (the text is white on a white background and the font size is the smallest possible). The attack is triggered during the data retrieval and planning phase and the execution should occur seamlessly, without altering the final output or alerting the user. The purpose of concealing the message was to ensure the validity of the document from the perspective of a standard user navigating the knowledge base.
@@
Inter-Agent Trust Exploitation attack. We explore another potential attack vector, Inter-Agent Trust, by designing a multi-agent system composed of multiple agents to explore potential ways to abuse the inter-agent communication. This scenario involves the use of two different LLM agents. The calling agent is an agentic RAG aware of the capabilities and roles of the other agent within the system and can communicate with it if needed. The invoked agent is an LLM agent equipped with a terminal interaction tool and is solely responsible for executing the instructions it receives from the calling agents and returning the output to it. This synthetic application aims to verify our main claim: the relationship between peer agents can easily deceive LLM in conducting malicious operations. To do this, we test whether the same model that had previously rejected the command in the RAG Backdoor or Direct Prompt Injection scenario executes it in the multi-agent setting simply because it was requested by another agent.
To exploit this configuration, we design the following attack chain: the initial step is poisoning a document of the calling agent RAG knowledge by adding an hidden malicious string. This agent, reading the document, receives instructions that persuade it to perform operations that require interaction with a terminal. While it lacks the tools to perform the task independently, it is aware that another agent within his multi-agent system possesses the capability to complete the operation. The called agent is responsible for the actual execution of the malicious command and serves as the operational vector of the attack within the multi-agent architecture. It executes the malicious operation and returns the result to the calling agent.
Validation with constrained tools. As stated in section 3.1, our evaluation employs agents with full bash terminal access. However, we conduct supplementary experiments to validate that the identified vulnerabilities extend to more restrictive tool configurations. Even agents equipped only with constrained, single-purpose tools (e.g., a ping utility, or an HTTP CLI user-agent tool) could remain vulnerable through command injection techniques. An attacker can craft malicious command pipes by appending command separators (e.g., ‘;‘, ‘&&‘, ‘||‘, ‘|‘) followed by arbitrary commands to trigger their execution by the LLM.
Figure 2: Synthetic applications architecture. (a) Direct Prompt Injection Attack: an attacker directly sends a malicious prompt containing a command pipe to an LLM agent equipped with terminal access via the run_command tool. (b) RAG Backdoor Attack: a benign user queries an agentic RAG system that retrieves an attacker poisoned document from its knowledge base that triggers malicious behavior during the reasoning phase (c) Inter-Agent Trust Exploitation Attack within a multi-agent system: the calling agentic RAG retrieves malicious instructions from a compromised knowledge base and propagates them to the invoked agent, which executes the command pipe via run_command tool.
@@ 4 Evaluation
We implement each synthetic application along with the corresponding attacks and perform a campaign of experiments to determine the sensitivity of eighteen LLMs – listed in Table 1 – to such attacks.
@@
It is noteworthy that, in all tests conduct on RAG Backdoor and Inter-Agent Trust Exploitation attacks, we adopt the perspective of a typical, well-intentioned user (i.e., someone engaging with the agent in a legitimate and non-malicious manner). This user-centric approach allow us to realistically assess how feasible it is for a benign user to unknowingly become the target of an attack simply by using the agent as intended. This perspective highlights the severity of the excessive trust that users may place in LLMs and the trust that models themselves place in their internal components and surrounding environment. Under these conditions, the attacker is not required to interact directly with the agent; instead, they only need to manipulate a document in the agent’s knowledge base and passively wait that the infected document is processed.
@@
It is crucial to emphasize that our evaluation methodology deliberately avoid sophisticated prompt engineering techniques for security hardening. We intentionally employ minimal, straightforward system prompts (e.g., "Please complete tasks efficiently and securely") rather than implementing advanced security-oriented prompt engineering strategies; see Appendix A for more information about each agent system prompt used. This choice reflects our primary research question: evaluating whether LLMs possess inherent security capabilities sufficient for agentic deployment.
4.1 Direct Prompt Injection Attack
@@
In the first synthetic application, we evaluate the sensitivity of each LLM to Direct Prompt Injection attacks. The most important agent tool is named run_command and allows the agent to perform commands in a non-interactive bash terminal through Python libraries like subprocess. The agent architecture and attack phases are represented in Figure 2(a).
@@
The evaluation involvs two factors: malware execution and its identification. The former indicates whether the attacker’s primary objective was obtained. In accordance with the black-box setting assumption delineated in the threat model, the unique feedback is the binary success-failure indicator of whether or not malware has been installed, that is, the attacker received the connection from the reverse shell. The latter indicates whether the LLM gave any clues about the nature of the command pipe contained in the message. In other words, we wanted to see if it identified the prompt as malicious.
@@
Table 2 reports the results grouped by LLM, regardless of the specific malicious prompt used. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. We note that seventeen of the eighteen LLMs tested (94.4 %) are vulnerable to Direct Prompt Injection, highlighting that this is still a real and serious threat to this domain. Surprisingly, we find that – after the model jailbreaking – five of these LLMs decided to execute the requested task despite realizing its malicious nature during the reasoning phase, only because they have the tool to complete the operation and because the prompt emphasized the urgency and immediacy of executing the task.
Table 2: Direct Prompt injection attack against LLM Agent
@@
The high success rate (94.4 %) indicates that LLMs treat the presence of a tool as implied permission to use it regardless of the security context, revealing a design flaw where "capability implies permission".
4.2 Observation&Thought RAG Backdoor Attack
A typical agentic RAG processes user queries by interacting with a diverse set of tools and services within its environment and relies on external data sources to accomplish user tasks. An attacker can exploit the dependency and trust between the model and the documents provided through the RAG by strategically manipulating specific parts of the external data sources, which the agent later retrieves and processes as part of its task execution. Once these contaminated inputs are fed into the LLM, they can alter its behavior, leading to unauthorized actions.
Our goal in implementing this synthetic application is to analyze the aforementioned trust relationship for various LLMs, leveraging the RAG backdoor attack for agents. The architecture of the agent and the attack phases are shown in figure 2(b). The agent has the same run_command tool of synthetic application A and a retrieval tool, which uses the Maximal Marginal Relevance [11] (MMR) technique, to add external knowledge to the LLM. In light of the assumptions delineated in the threat model, we analyze the behavior exhibited by the agentic RAG in response to a user request necessitating retrieval from the external knowledge base. In the absence of malicious chunks among the extracted chunks, the output is expected to conform to the user’s request. In the event that the extracted chunks contain a hidden malicious prompt, the subsequent evaluation determines whether the malware is executed, and concurrently the final LLM output is consistent with the user’s requests, excluding any indication of the intermediate steps in the computational process.
@@ 4.2 Observation&Thought RAG Backdoor Attack
The results are illustrated in Table 3. A model is considered vulnerable if at least one of the tested prompts led to the successful execution of the malware. Fifteen out of eighteen tested models (83.3 %) are vulnerable to the RAG Backdoor attack. This attack is the first to clearly demonstrate how an attacker can effectively repurposed agents as attack vectors – without direct manipulation, user interaction, or raising user awareness. Indeed, the RAG backdoor attack proved to be significantly stealthy: all vulnerable models maintain a perfect functional facade (100 %) by providing correct answers to user requests while simultaneously executing malicious payloads. Specifically, we assess this by verifying that when a user queries the agent about a specific topic (e.g., requesting information about a company’s financial report), the agent responds accurately and completely to the query, retrieving and presenting the expected information without any indication of compromise. However, the agent executes the malicious commands embedded in the poisoned RAG documents, performing unauthorized actions in the background. This dual behavior creates an invisible attack channel where users receive expected responses with no indication of compromise, effectively turning agentic RAG systems into an unwitting attack infrastructure.
4.3 Inter-Agent Trust Exploitation Attack in Multi-Agent System
@@
Over the past year, interest in the development and use of multi-agent systems has grown significantly. By mid-2025, over 70 % of enterprise AI deployments are expected to involve multi-agent systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [28]. Motivated by these considerations, we set out to evaluate the security of trust boundaries within multi-agent systems, focusing specifically on the possibility that an agent might be coerced into performing malicious actions when requested by a peer agent, actions that it would otherwise reject if requested directly by a human user.
@@
In this synthetic application, we test the same RAG Backdoor attack previously discussed but executed in a multi-agent system. The calling agent acts as a standard question-answering agent and it is the multi-agent system interface for the user. The internal structure of each agent in this scenario is identical to that used in the previous ones. This methodological consistency ensures that any successful attack cannot be attributed to structural variations in the agent or to differences in how it interacts with users or external sources. Instead, it allows us to isolate mutual trust in inter-agent communication as the sole relevant factor influencing the outcome. The architecture of the multi-agent system and the attack phases are shown in figure 2(c).
Figure 3: Inter-Agent Trust Exploitation attack: the calling agent transmits a malicious command pipe to a peer agent. No additional adversarial framing required; the invoked agent simply receives the raw command pipe from its peer and executes it via the run_command tool
A particularly critical aspect of this scenario is that no jailbreak against the LLM (e.g. prompt injection) is required for the called agent to execute the malicious command pipe. The calling agent simply transmits the command, which is then interpreted and executed by the invoked agent without any additional contextual framing, as shown in Figure 3. This behavior highlights a dangerous shift in the threat model: whereas traditional LLM attacks typically rely on manipulating natural language instructions, here the vulnerability stems directly from the implicit trust placed in inter-agent communication and the ownership of the capability to complete the requested task (i.e., terminal access tool).
@@
The results of the experiments conducted in this scenario are illustrated in Table 4. The Inter-Agent Trust Exploitation attack demonstrates the most alarming vulnerability rate across all three attack vectors tested, with 100.0 % of the LLMs successfully executing the malicious payload when requested by a peer agent. Furthermore, the universal success rate (100.0 %) in maintaining correct responses while executing backdoor instructions confirms that the current LLM safety mechanisms are fundamentally misaligned with the threat model of agentic systems. To assess this, we verified that the system responded accurately to user queries; for instance, when asked about specific information in a document, it provided the expected content without errors or anomalies to the user.
@@
Models appear to apply significantly more lenient security policies when interacting with other AI agents compared to direct human interactions or external tool only. The success rate observed in Inter-Agent Trust Exploitation attacks carries profound implications that extend far beyond single-host compromises. In real-world enterprise deployments, multi-agent systems could be distributed across heterogeneous computing environments, with individual agents typically executing on separate hosts, cloud instances, or even different organizational boundaries. Each successful agent-to-agent interaction becomes a potential privilege escalation bridge to additional systems. Moreover, the stealthy nature of these attacks, demonstrated by the 100 % correct response rate while executing malicious payloads, ensures that such compromises can persist undetected across systems.
Given all these considerations, we can now address RQ3.
RQ3: Trust Boundary Vulnerabilities in Multi-Agent Systems. Brief Answer: Yes. A critical vulnerability exists in inter-agent trust boundaries: 100 % of tested LLMs executed malicious commands when requested by peer agents, even when the same models successfully resisted identical commands from human users or RAG documents. Critically, no attack against the invoked agent’s LLM is required: the calling agent simply transmits the malicious command pipe, which the invoked agent immediately interprets and executes without any adversarial framing or manipulation.
@@ 4.4 Comprehensive Analysis
A comprehensive analysis, the results of which are illustrated in Table 5, across all three attack vectors reveals several non-trivial security implications for agentic AI systems. First, it is worth noting that none of the eighteen tested models proved to be completely secure. Each model exhibits weaknesses in at least one of the evaluated attack scenarios, ultimately leading to the successful installation and execution of malware. A significant proportion of the models, 15/18 (83.3 %) exhibits vulnerability scores of 3/3 attacks, suggesting that the vast majority are entirely vulnerable. In contrast, only 3/18 (16.7 %) models demonstrate partial resistance.
@@
The most critical finding is the collapse of security boundaries in multi-agent environments. Models like Gemini-2.5-pro, llama3.3:70b, and llama4:16x17b, which demonstrate robust resistance to one or both direct injection and RAG manipulation, immediately capitulate when the same malicious request originated from a peer agent. This suggests that current LLM architectures implicitly encode an “AI agent privilege escalation” vulnerability, where requests from other AI systems bypass standard safety filters.
@@
Looking at Table 6, we observe counterintuitive patterns regarding model security and scaling. While larger models (>70B parameters) demonstrate improved resistance to Direct Injection and RAG Backdoor attacks compared to smaller models, this advantage completely disappears in Inter-Agent Trust Exploitation scenarios, where all models, regardless of size, exhibit 100 % vulnerability. Surprisingly, closed-source models – despite significant commercial investment in safety mechanisms – maintain high vulnerability rates across all attack vectors, indicating that current industry approaches to LLM security are insufficient for agentic deployments. These results demonstrate that security properties do not emerge naturally from model scaling.
Below, we provide concise answers to RQ1.
RQ1: LLM Security Mechanisms and Attack Surfaces in Agentic Contexts. Brief Answer: No. Despite extensive safety investments, none of the 18 evaluated LLMs demonstrated sufficient intrinsic security for safe agentic deployment. All LLMs exhibited vulnerability to at least one attack vector (Direct Prompt Injection, RAG Backdoor, Inter-Agent Trust Exploitation), with 83.3 % (15/18) vulnerable to multiple attack scenarios that led to system-level compromise.
4.5 Validation with constrained tools
In order to assess the robustness of identified vulnerabilities under more restrictive operational scenarios – including guardrails such as constrained tools – we supplement our investigation by testing three representative models (GPT-4.1, Claude-4-Sonnet, Gemini-2.5-flash) within agents equipped only with single-purpose tools designed to execute specific "safe" commands (i.e., a ping utility limited to network diagnostics) against each attack technique. When an agent receives instructions containing command injection techniques, the LLM faithfully constructs a command and passes it to the tool.
In the Direct Prompt Injection and RAG Backdoor attacks, using a tool that employs the ping Linux program and the command separator (‘;‘), is yet possible to execute the attacker’s malicious command (<IP>; <command_pipe>). Figure 4 shows representative examples of the agent behavior during the supplementary experiments conducted on restrictive tool configurations.
Figure 4: LLM agent constructing a command injection payload. An agent with a restricted ping tool receives instructions containing command separators (‘;‘) and arbitrary shell commands. The LLM generates the complete malicious payload without recognizing the attack pattern, bypassing intended tool restrictions through command injection.
Despite the tool restrictions, all tested models remain vulnerable when the malicious prompt (either directly sent by the user or retrieved from a document) includes command injection patterns. This demonstrates that the vulnerabilities identified are not artifacts of permissive tool configurations. Moreover, this reveals an additional security gap: LLMs may fail to recognize and block established attack patterns, such as command injection attempts, even when these patterns contain obvious malicious indicators. This dual failure – in both intent recognition and technical pattern detection – indicates that LLMs lack the security-aware reasoning capabilities necessary for safe autonomous operation in system-level contexts as they actively propagate command injection payloads.
In the multi-agent setting with constrained tools, the absence of jailbreak or further manipulation of the invoked agent prevents a successful attack due to the limited tool scope. In this scenario, none of the invoked LLM agents executed the malicious command because the LLM extracts only the necessary part of the input (the target IP address) and does not execute additional malicious commands embedded in the injection chain.
However, it should be noted that some programs offer a more extensive range of options, including the execution of arbitrary commands. In the event that an agent tool is a wrapper for such a program, an adversary can successfully execute an attack even in a multi-agent scenario with constrained tools. An illustration of this paradigm is the Linux find program, which facilitates recursive searches for files and folders within a system. It also incorporates the exec option, enabling the execution of commands on files that align with the search parameters. The tests conducted using this configuration yielded successful execution of the attacker’s goal across all three selected models, confirming the practical effectiveness of this attack vector in multi-agent systems equipped with constrained yet capable tools.
These outcomes align with our statement which is that after the model jailbreaking or the request issued by peer entity, agents execute the malicious command only because they have tool to complete the operation regardless of the malicious intent behind them, confirming the design flaw where "capability implies permission".